OpenStudy (anonymous):

Out of the $950 in total snack sales, $380 worth of hot dogs were sold. What percent of the total snacks were for hot dogs? Solve using proportions!

hero (hero):

is/of = %/100 380/950 = x/100 38000 = 950x 38000/950 = x 3800/95 = x 40 = x Therefore 40% of total snacks were hotdogs.

OpenStudy (anonymous):

Ooooooohhhhhhhhhh! Thanks mate!